OpenStudy (anonymous):

First estimate the answer. Then subtract to find the exact answer. 12 2/9 - 7 7/15 = but i have to round each first

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what does 12 2/9 round to?

OpenStudy (anonymous):

11right

jimthompson5910 (jim_thompson5910):

close, but no

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

now you're getting further away

OpenStudy (anonymous):

its 12 sorry

jimthompson5910 (jim_thompson5910):

7 7/15 rounds to what?

OpenStudy (anonymous):

stays the same

jimthompson5910 (jim_thompson5910):

so it rounds to 7?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 12 2/9 - 7 7/15 turns into 12 - 7 = 5

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

that's the estimated answer

jimthompson5910 (jim_thompson5910):

any ideas on how to get the exact answer?

OpenStudy (anonymous):

yeah it askes for that too and no i dont know how to do this math

jimthompson5910 (jim_thompson5910):

did you see how I did the last problem?

OpenStudy (anonymous):

yes but i get confused im not good with with multiplication

jimthompson5910 (jim_thompson5910):

do you know how to find the LCD?

OpenStudy (anonymous):

yes common factors right of 9 and 15

jimthompson5910 (jim_thompson5910):

ok what is the GCF of 9 and 15?

OpenStudy (anonymous):

45

jimthompson5910 (jim_thompson5910):

that's the LCM, but that's what we wanted anyway

jimthompson5910 (jim_thompson5910):

so we need to get each denominator equal to the LCM

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

So we do this \[\large 12 \frac{2}{9} - 7 \frac{7}{15}\] \[\large 12 \frac{2*5}{9*5} - 7 \frac{7*3}{15*3}\] \[\large 12 \frac{10}{45} - 7 \frac{21}{45}\]

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

we subtract it right

jimthompson5910 (jim_thompson5910):

not yet

jimthompson5910 (jim_thompson5910):

we need to convert over to improper fractions

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so what do you get when you do so?

OpenStudy (anonymous):

dont we multiply the whole number with the denominator and then add the numerator

jimthompson5910 (jim_thompson5910):

good, you are correct

jimthompson5910 (jim_thompson5910):

so do that for each mixed number

OpenStudy (anonymous):

550/45- 336/45 right?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now subtract

jimthompson5910 (jim_thompson5910):

this is now possible because we have 2 fractions (with no whole parts) and the denominators are the same

OpenStudy (anonymous):

214/45 right?

jimthompson5910 (jim_thompson5910):

now convert that back to a mixed number

OpenStudy (anonymous):

ok i dont know how to do that dont i divide 45 by 214

jimthompson5910 (jim_thompson5910):

since 214/45 = 4.75555555555556 (use a calculator) this means the quotient is 4 take the decimal part 0.75555555555556 and multiply it by 45 to get 0.75555555555556*45 = 34.0000000000002 which is practically 34 so the remainder is 34 ------------------------------------------------------- 214/45 = 4 remainder 34 so... \[\large \frac{214}{45} = 4\frac{34}{45}\]

jimthompson5910 (jim_thompson5910):

and that means \[\large 12 \frac{2}{9} - 7 \frac{7}{15} = 4\frac{34}{45}\]

OpenStudy (anonymous):

oh okay thank you can you help me with two morte

jimthompson5910 (jim_thompson5910):

yeah sure, but I have to do something real quick, so I'll brb